OpenStudy (anonymous):

what is 17 of 58? 1.5 is 45% of what number? what is 12.5% of 34.50? what percent is 22 of 222?

OpenStudy (anonymous):

can anyone help

OpenStudy (anonymous):

Okay, so there is one word that we see in each of the four lines and that is "of". In math language, "of" often means multiply as in the first question: Q1: 17 * 58 = ? Also "is" means equal to and "what number" is our unknown variable "y" Q2: 1.5 = (0.45)*(y) {so use algebra to solve for y.} Q3: ? = (0.125)* (34.50) Q4: This last one is a bit tricky, and you need to divide the first number by the second number.